OpenStudy (anonymous):

ASAP PLEASE ; What is the 20th term of the arithmetic sequence 3, –3, –9, … ?

OpenStudy (anonymous):

please help! i will give medal

OpenStudy (anonymous):

111 This is an Arithimatic progression series. Hence the formula for the 20 term is t(20)= -3 + (20-1)*6 = 111

OpenStudy (anonymous):

wouldnt it be -111?

OpenStudy (anonymous):

ok i thought the series in +ve. the answer is: t(20) = -3 + (20-1)* -6 = -117

OpenStudy (anonymous):

If the first term is 3 and the second is -3: ... 3, -3, -9, -15, -21, -27, -33, -39, -45, -51, -57, -63, -69, -75, -81, -87, -93, -99, -105, -111.